× | products . Data Briefs-k × | O etext ps://berkeley.courseload.com/#/content-43581/address/133 xPage 85, Fundamentals of C Sylilabus for CO X ESTIONS AND PROBLEMS Calculating Liquidity Ratios [LO2] SDJ, Inc., has net working capital of $1,920, current liabilities of $4,380, and inventory of $3,750. What is the current ratio? What is the quick ratio? Calculating Profitability Ratios LO2] Shelton, Inc., has sales of $17.5 million, total assets of $13.1 million, and total debt of $5.7 million. If the profit margin is 6 percent, what is net income? What is ROA? What is ROE? Calculating the Average Collection Period [LO2] Aguilera Corp. has a current accounts receivable balance of $438,720. Credit sales for the year just ended were $5,173,820. What is the receivables turnover? The days sales in receivables? How long did it take on average for credit customers to pay off their accounts during the past year? Calculating Inventory Turnover [LO2] The Green Corporation has ending inven- tory of $417,381, and cost of goods sold for the year just ended was $4,682,715. What is the inventory turnover? The days sales in inventory? How long on average did a unit of inventory sit on the shelf before it was sold? Calculating Leverage Ratios [LO2] Levine, Inc., has a total debt ratio of .53. What is its debt-equity ratio? What is its equity multiplier? Calculating Market Value Ratios [LO2] Makers Corp. had additions to retained earnings for the year just ended of $395,000. The firm paid out $195,000 in cash dividends, and it has ending total equity of $5.3 million. If the company currently to search
Answer #1

1. Calculating Liquidity Ratios

Current Ratio = (Net Working Capital + Current Liabilities)/Current Liabilities

Current Ratio = (1920+4380)/4380

Current Ratio = 1.438

Quick Ratio = (Current Assets-Inventory)/Current Liabilities

Quick Ratio = (1920+4380-3750)/4380

Quick Ratio = 0.582

2. Calculating Profitability Ratios

Net Profit = Sales * Profit Margin

Net Profit = 17.5 * 0.06

Net Profit = $1.05 million

ROA = Net Profit/Total Assets

ROA = 1.05/13.1

ROA = 8.01%

ROE = Net Profit.(Assets - Debt)

ROE = 1.05/(13.1-5.7)

ROE = 14.12%

3. Calculating Average Collection Period

Receivables Turnover = Credit Sales/Account Receivables

Receivables Turnover = 5173820/438720

Receivables Turnover = 11.79

Days Sales Receivables = 365/Receivables Turnover

Days Sales Receivables = 365/11.79

Days Sales Receivables = 30.95 days

4. Calculating Inventory Turnover

Inventory Turnover = COGS/Invnetory

Inventory Turnover = 4682715/417381

Inventory Turnover = 11.22

Day's Sales Inventory = 365/Inventory turnover

Day's Sales Inventory = 32.53 days